# Aspects of holographic Langevin diffusion in the presence of anisotropic magnetic field

### v1

@@ Abstract
We revisit the holographic Langevin diffusion coefficients of a heavy quark, when travelling through a strongly coupled anisotropic plasma in the presence of magnetic field $\mathcal{B}$. The Langevin diffusion coefficients are calculated within the membrane paradigm in the magnetic branes model which has been extensively studied to investigate the magnetic effects on various observables in strongly coupled QCD scenarios by holography. In addition to confirming some conventional conclusions, we also find several new interesting features among the five Langevin diffusion coefficients in the magnetic anisotropic plasma. It is observed that the transverse Langevin diffusion coefficients depend more on the direction of motion rather than the directions of momentum diffusion at the ultra-fast limit, while one would find an opposite conclusion when the moving speed is sufficiently low. For the Longitudinal Langevin diffusion coefficient, we find that motion perpendicular to $\mathcal{B}$ affects the Langevin coefficients stronger at any fixed velocity. We should also emphasize that all five Langevin coefficients are becoming larger with increasing velocity. We find that the universal relation $\kappa^{\parallel}>\kappa^{\perp}$ in the isotropic background, is broken in a different new case that a quark moving paralleled to $\mathcal{B}$. This is one more particular example where the violation of the universal relation occurs for the anisotropic background. Further, we find the critical velocity of the violation will become larger with increasing $\mathcal{B}$.
@@ I Introduction
A strong magnetic field is produced in heavy ion collisions (HICs) experiments at the Relativistic Heavy Ion Collider (RHIC) around $e\mathcal{B}\sim 0.02GeV^{2}$ and the Large Hadron Collider (LHC) up to $e\mathcal{B}\sim 0.3GeV^{2}$ Skokov et al. (2009); Deng and Huang (2012); Bloczynski et al. (2013); Tuchin (2013); Bali et al. (2012), which has attracted much attention to study the influences on properties of the quark gluon plasma (QGP) from the magnetic field. It’s also known the existence of even stronger magnetic fields in the interior of the dense neutron stars Duncan and Thompson (1992) and early stages of the Universe Vachaspati (1991); Grasso and Rubinstein (2001). In weakly coupled quantum chromodynamics (QCD) theory, the influences from the strong magnetic field $\mathcal{B}$ to dynamic heavy quark has been studied at the Lowest-Landau-Level approximation, which for a thermal medium suggests the regime $e\mathcal{B}\gg T^{2}$ Fukushima et al. (2016); Kurian et al. (2019); Sadofyev and Yin (2016); Zhang et al. (2021); Bandyopadhyay et al. (2022).
@@
The AdS/CFT correspondence Maldacena (1998); Witten (1998a); Gubser et al. (1998) is one of the very promising approaches to deal with these problems in QCD at strong couple scenario where can’t be perturbatively handled properly Witten (1998b); Aharony et al. (2000). On gravity side, the heavy quark diffusion in a strong coupled plasma can be understood as the fluctuation correlations of the trailing string. The studying of the stochastic nature of a heavy quark in a holography was proposed by Gubser (2008); Casalderrey-Solana and Teaney (2007). Soon the stochastic motion is formulated as a Langevin process de Boer et al. (2009); Son and Teaney (2009). Since the heavy quarks in HICs experiments are relativistic in many cases, the relativistic Langevin evolution are studied in Giecold et al. (2009) and in non-conformal frameworks in Gursoy et al. (2010); Kiritsis et al. (2012) for the multiple scales of QCD. The effect of the magnetic field directly on the heavy quark moving was studied in Matsuo et al. (2006); Kiritsis and Pavlopoulos (2012), which ignores the effect of the magnetic field on the plasma. It is studied that the heavy quark moving in magnetized plasma with $\gamma M\leq\sqrt{\mathcal{B}}$ ($\gamma$ is Lorentz factor and $M$ is heavy quark mass), at strong magnetic field limit $e\mathcal{B}\gg T^{2}$ Mamo (2016); Li et al. (2016). Other aspects of influences from magnetic field to heavy quark with holography Zhu et al. (2022); Zhang and Ma (2018); Zhang et al. (2018); Zhang (2019); Zhu et al. (2019, 2020); Zhou et al. (2020); Chen et al. (2018); He et al. (2013); Dudal and Mertens (2018); Giataganas (2018); Iwasaki et al. (2021); Chakrabortty et al. (2014); Rougemont (2020); Fischler et al. (2014); Rougemont et al. (2015); Bu et al. (2022); Bu and Zhang (2021).
@@
An universal inequality relation, $\kappa_{\parallel}>\kappa_{\perp}$, has been found in the generic isotropic background with AdS/CFT correspondence Gursoy et al. (2010); Pourhassan et al. (2017). A way to violate this relation is to consider the Brownian motion in anisotropic background. And the fact that the transverse LGV-coefficient may be larger than a specific longitudinal LGV-coefficient, is found in both JW and MT spatial anisotropic background in Giataganas and Soltanpanahi (2014a). Due to the rapid expansion along the beam direction, the QGP experiences an anisotropic phase both in momentum and coordinate space before the system reaches the isotropic phase in a short period of time. And the strong magnetic field is another important source of anisotropy, and it is interesting to study its influence on heavy quark diffusion at HICs. To be specific we want to investigate how the anisotropy induced by an uniform magnetic field affects the both longitudinal and transverse LGV-coefficients.
In this work, we will study LGV-coefficients with the magnetic branes model to present our findings. Besides the confirmation of some conventional conclusions found before, we will conduct an in-depth investigation about heavy quark diffusion with LGV-coefficients in the presence of a uniform magnetic field with holography. We also want to study the heavy quark diffusion in an anisotropic background with a uniform magnetic field $\mathcal{B}$. The relations between LGV-coefficients in an anisotropic background also is an important motivation for this work. Also, we hope our study could shed light on to transport and hydrodynamic Monte Carlo simulation of HICs.
@@ II The setups
then he full bulk metric (2) after these rescalings reads
$\begin{split}ds^{2}=&-\tilde{U}(\tilde{r})d\tilde{t}^{2}\\ &+\frac{e^{2\tilde{V}(\tilde{r})}}{v(b)}(d\tilde{x}^{2}+d\tilde{y}^{2})+\frac{e^{2\tilde{W}(\tilde{r})}}{w(b)}d\tilde{z}^{2}+\frac{d\tilde{r}^{2}}{\tilde{U}(\tilde{r})}\\ F=&\frac{b}{v(b)}d\tilde{x}\wedge d\tilde{y}.\end{split}$ (16)
@@
Next, we compute the LGV-coefficients of a heavy quark in a uniform magnetic field induced anisotropy medium by following the author of Gursoy et al. (2010); Giataganas and Soltanpanahi (2014a); Finazzo et al. (2016). We consider a general diagonal metric in radial coordinates,
@@
Following the usual procedure, the effective temperature reads
$T^{2}_{ws}$ $=\frac{1}{16\pi^{2}}\absolutevalue{\frac{g^{{}^{\prime}2}_{tt}-v^{4}g^{{}^{\prime}2}_{pp}}{g_{tt}g_{pp}}}\bigg{|}_{r=r_{c}}$ (26)
$=\frac{1}{16\pi^{2}}\absolutevalue{\frac{1}{g_{tt}g_{rr}}(g_{tt}g_{pp})^{\prime}(\frac{g_{tt}}{g_{pp}})^{\prime}}\bigg{|}_{r=r_{c}}.$
@@
[…]
Taking advantage of the membrane paradigm Iqbal and Liu (2009), one can directly reads LGV-coefficients without solving motion equation,
@@
$T^{\parallel}_{ws}$ $=\frac{1}{4\pi}\left[\absolutevalue{\left(\frac{e^{2\tilde{W}(\tilde{r})}\tilde{U}(\tilde{r})}{w(b)}\right)^{\prime}\left(\frac{w(b)\tilde{U}(\tilde{r})}{e^{2\tilde{W}(\tilde{r})}}\right)^{\prime}}\right]^{\frac{1}{2}}\bigg{|}_{r=r_{c}}.$ (35)
@@
$T^{\perp}_{ws}$ $=$ $=\frac{1}{4\pi}\left[\absolutevalue{\left(\frac{e^{2\tilde{V}(\tilde{r})}\tilde{U}(\tilde{r})}{v(b)}\right)^{\prime}\left(\frac{v(b)\tilde{U}(\tilde{r})}{e^{2\tilde{V}(\tilde{r})}}\right)^{\prime}}\right]^{\frac{1}{2}}\bigg{|}_{r=r_{c}}.$ (37)
@@ III.1 Directions dependence
Initially, in this subsection, we will examine two extreme cases to study the relations between different LGV-coefficients in all scale with $\mathcal{B}/T^{2}$, for which is the only dimensionless parameter in this model. The numerical results of transverse LGV-coefficients $\kappa_{\perp}$ and longitudinal LGV-coefficients $\kappa_{\parallel}$ as a function of magnetic field $B/T^{2}$ are displayed in Fig. 2 and Fig.3, normalized by the conform limit given in (40). Fig. 2 show transverse LGV-coefficients $\kappa^{v\parallel B}_{\perp}$, $\kappa^{v\perp B}_{(\perp,\parallel)}$ and $\kappa^{v\perp B}_{(\perp,\perp)}$, while Fig. 3 show longitudinal LGV-coefficients $\kappa^{v\parallel B}_{\parallel}$ and $\kappa^{v\perp B}_{\parallel}$.
@@
When moving at the ultra fast limit, as shown in plot (a) in Fig. 2, the transverse LGV-coefficients of quark $\kappa_{\perp}$ has a relation reads $\kappa_{\perp}^{v\parallel B}<\left(\kappa^{v\perp B}_{(\perp,\parallel)}\sim\kappa^{v\perp B}_{(\perp,\perp)}\right)$, which means the transverse LGV-coefficients of quark mostly depending on the direction of moving but not the direction of transverse momentum diffusion when the moving velocity approaching to light speed. It also means that the anisotropy in plane $xoz$ contributes little to the strength of transverse LGV-coefficients when moving in ultra-fast limit. But when the moving velocity is sufficiently small, $v=0.1$ is taken for example, one will find an opposite conclusion that the transverse LGV-coefficients of quark depend more on the direction of transverse momentum diffusion rather than the direction of moving.
As for the longitudinal LGV-coefficients shown in Fig. 3, we find that $\kappa_{\parallel}^{v\perp B}<\kappa_{\parallel}^{v\parallel B}$ at both ultra fast limit and a fixed sufficiently small speed. And moving perpendicular to $\mathcal{B}$ direction affects the longitudinal LGV-coefficients stronger compared to motion along $\mathcal{B}$ direction. So we conclude that the transverse LGV-coefficients $\kappa_{\perp}$ may be more sensitive to the anisotropic magnetic background rather than the Longitudinal LGV-coefficients $\kappa_{\parallel}$.
By comparing transverse LGV-coefficients in Fig. 2 and the longitudinal LGV-coefficients in Fig. 3, we find that the transverse LGV-coefficients may be more sensitive to the anisotropic magnetic background rather than the Longitudinal LGV-coefficients. Finally, the LGV-coefficient always enhances with increasing magnetic field $\mathcal{B}$ in all cases for both transverse and longitudinal LGV-coefficients.
@@
III.2 Dynamic properties
We further examine the impact of moving velocity $v$ on both transverse and longitudinal LGV-coefficients. In Fig. 4, we plot the LGV-coefficients as a function of quark moving velocity $v$, accompanied by the corresponding results in conformal limit at (40). It demonstrates that the increasing moving velocity $v$ enhances the LGV-coefficients in all cases we mentioned in this anisotropic magnetic background, which is very similar to the results in the conformal limit. This conclusion is also similar to the outcomes in spatial anisotropic case in Giataganas and Soltanpanahi (2014b).
Unlike the two extreme cases with fixed speeds in III.1, we can study the same question with varying moving speeds. It is found that, at a sufficiently small moving speed, the momentum broadening of a heavy quark depends more on the direction of momentum diffusion rather than the direction of moving with both $\kappa^{v\perp B}_{(\perp,\parallel)}<\left(\kappa^{v\perp B}_{(\parallel,\perp)}\sim\kappa^{v\parallel B}_{\perp}\right)$ in the upper panel and $\kappa^{v\parallel B}_{\parallel}<\kappa^{v\perp B}_{\parallel}$ in the lower panel of Fig. (4).

### v2

@@ Abstract
We study the holographic Langevin diffusion coefficients of a heavy quark, when traveling through a strongly coupled anisotropic plasma in the presence of magnetic field $\mathcal{B}$. While previous studies have characterized Langevin diffusion coefficients in the magnetic branes model, our analysis uncovers several new features among the five coefficients in this magnetic anisotropic plasma, advancing the understanding of heavy quark dynamics in such environments. In particular, we clarified how motion velocities shape momentum broadening and its directional dependence. It is observed that the transverse Langevin diffusion coefficients depend more on the direction of motion rather than the directions of momentum diffusion at the ultra-fast limit, while an opposite conclusion is found when the moving speed is sufficiently low. For the Longitudinal Langevin diffusion coefficient, we find that motion perpendicular to $\mathcal{B}$ affects the Langevin coefficients stronger at any fixed velocity. We should also emphasize that all five Langevin coefficients are becoming larger with increasing velocity. We find that the universal relation $\kappa^{\parallel}>\kappa^{\perp}$ in the isotropic background, is broken in a different new case that a quark moving paralleled to $\mathcal{B}$. This is one more particular example where the violation of the universal relation occurs for the anisotropic background. Further, we find the critical velocity of the violation will become larger with increasing $\mathcal{B}$.
@@ I Introduction
The heavy-ion collisions (HIC) experiments at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) are believed to create almost the most perfect fluid Quark Gluon Plasma (QGP) Arsene et al. (2005); Adcox et al. (2005); Back et al. (2005); Adams et al. (2005). This provides a novel window for studying the physics of Quantum Chromodynamics (QCD) at a strongly coupled regime. Since the properties of a strongly coupled system cannot be reliably calculated directly by perturbative techniques, one has to resort to some nonperturbative approaches to overcome the challenges.
@@
The AdS/CFT correspondence Maldacena (1998); Witten (1998a); Gubser et al. (1998) is one of the very promising approaches to deal with these problems in QCD at strong couple scenario where can’t be perturbatively handled properly Witten (1998b); Aharony et al. (2000). On gravity side, the heavy quark diffusion in a strong coupled plasma can be understood as the fluctuation correlations of the trailing string. The studying of the stochastic nature of a heavy quark in a holography was proposed by Gubser (2008); Casalderrey-Solana and Teaney (2007). Soon the stochastic motion is formulated as a Langevin process de Boer et al. (2009); Son and Teaney (2009). Since the heavy quarks in HIC experiments are relativistic in many cases, the relativistic Langevin evolution are studied in Giecold et al. (2009); Casalderrey-Solana et al. (2009) as well as in non-conformal frameworks in Gursoy et al. (2010); Kiritsis et al. (2012, 2014) towards the multiple scales of QCD. Recently, the contribution of nonlinear terms has been considered in Chakrabarty et al. (2020); Chakrabarty and P. M. (2021); Bu and Zhang (2021); Bu et al. (2022) using a new prescription for Schwinger-Keldysh correlators proposed in Glorioso et al. (2018). The effect of the magnetic field directly on the heavy quark moving was studied in Matsuo et al. (2006); Kiritsis and Pavlopoulos (2012), which ignores the effect of the magnetic field on the plasma. It is studied that the heavy quark moving in magnetized plasma with $\gamma M\leq\sqrt{\mathcal{B}}$ ($\gamma$ is Lorentz factor and $M$ is heavy quark mass), at strong magnetic field limit $e\mathcal{B}\gg T^{2}$ Mamo (2016); Li et al. (2016). Other aspects of the influence of the magnetic field on heavy quarks using holography Zhu et al. (2022); Zhang and Ma (2018); Zhang et al. (2018); Zhang (2019); Zhu et al. (2019, 2020); Zhou et al. (2020); Chen et al. (2018); He et al. (2013); Dudal and Mertens (2018); Giataganas (2018); Iwasaki et al. (2021); Chakrabortty et al. (2014); Rougemont (2020); Fischler et al. (2014); Rougemont et al. (2015); Pourhassan et al. (2017); Chakrabortty et al. (2014); Fischler et al. (2012); Tong and Wong (2013) and references therein.
@@
An universal inequality relation, $\kappa_{\parallel}>\kappa_{\perp}$, has been found in the generic isotropic background with AdS/CFT correspondence Giataganas and Soltanpanahi (2014a), based on the developments of the theory chosen in Gursoy et al. (2010). A way to violate this relation is to consider the Brownian motion in anisotropic background. And the fact that the transverse Langevin diffusion coefficients (LGV-coefficient) may be larger than a specific longitudinal LGV-coefficient, is found in both Janik-Witaszczyk Janik and Witaszczyk (2008) and Mateos-Trancanelli Mateos and Trancanelli (2011a, b) spatial anisotropic background in Giataganas and Soltanpanahi (2014b). Due to the rapid expansion along the beam direction, the QGP experiences an anisotropic phase both in momentum and coordinate space before the system reaches the isotropic phase in a short period of time. And the strong magnetic field is another important source of anisotropy, and it is interesting to study its influence on heavy quark diffusion at HIC. To be specific we want to investigate how the anisotropy induced by an uniform magnetic field affects the both longitudinal and transverse LGV-coefficients.
A strong magnetic field is produced in HIC experiments at RHIC around $e\mathcal{B}\sim 0.02GeV^{2}$ and LHC up to $e\mathcal{B}\sim 0.3GeV^{2}$ Skokov et al. (2009); Deng and Huang (2012); Bloczynski et al. (2013); Tuchin (2013); Bali et al. (2012), which has attracted much attention to study the influences on properties of QGP from the magnetic field. It’s also known the existence of even stronger magnetic fields in the interior of the dense neutron stars Duncan and Thompson (1992) and early stages of the Universe Vachaspati (1991); Grasso and Rubinstein (2001). In weakly coupled quantum chromodynamics (QCD) theory, the influences from the strong magnetic field $\mathcal{B}$ to dynamic heavy quark has been studied at the Lowest-Landau-Level approximation, which for a thermal medium suggests the regime $e\mathcal{B}\gg T^{2}$ Fukushima et al. (2016); Kurian et al. (2019); Sadofyev and Yin (2016); Zhang et al. (2021); Bandyopadhyay et al. (2022). The top-down magnetic branes model is dual to a deformed strong coupled SYM theory with a constant magnetic field D’Hoker and Kraus (2009, 2013, 2010). Except for the top-down approach to modeling QCD for example D’Hoker and Kraus (2009, 2013, 2010); Witten (1998c); Hata et al. (2007), the bottom-up method is also very popular in this area and has yielded many extraordinary results Erlich et al. (2005); Da Rold and Pomarol (2005); Karch et al. (2006); Gubser and Nellore (2008); Gubser et al. (2008); Gursoy et al. (2009, 2011); Critelli et al. (2017); Cai et al. (2012); Giataganas et al. (2018). Although these models may not be exactly the same as the real QCD, it is expected that this kind of effort could help to reveal some intrinsic features of the plasma.
In this work, we choose to study LGV-coefficients within the widely studied magnetic branes model D’Hoker and Kraus (2009, 2013, 2010) to present our findings. The authors of Finazzo et al. (2016) first computed LGV-coefficients for the magnetic branes model and discovered several interesting properties of these coefficients. In addition to confirming some conventional conclusions previously reported in Finazzo et al. (2016), we conduct an in-depth investigation of heavy quark diffusion using LGV-coefficients in the presence of a uniform magnetic field, employing the same strategies. We also aim to study heavy quark diffusion in an anisotropic background induced by a uniform magnetic field $\mathcal{B}$, and the relationships between LGV-coefficients in an anisotropic background are another important motivation for this work. Furthermore, we hope that our study can provide insights into transport simulations of hard probes in HIC.
@@ II The setups
then he full bulk metric (2) after these rescaling reads
[…]
@@
Next, we compute the LGV-coefficients of a heavy quark in a uniform magnetic field induced anisotropy medium mainly by following the author of Giataganas and Soltanpanahi (2014a) and also Finazzo et al. (2016); Gubser (2008); Gursoy et al. (2010). We consider a general diagonal metric in radial coordinates,
@@
As derived by the authors of Giataganas and Soltanpanahi (2014a, b), the effective temperature reads,
[…]
@@
$S_{f}=-\frac{1}{2\pi\alpha^{\prime}}\int d\tau d\sigma\frac{H^{\alpha\beta}}{2}\\ \times(N[r]\partial_{\alpha}\delta x_{p}\partial_{\beta}\delta x_{p}+\sum g_{ii}\partial_{\alpha}\delta x_{i}\partial_{\beta}\delta x_{i}),\\ N(r)\equiv\frac{g_{tt}g_{pp}+C^{2}}{g_{tt}+g_{pp}v^{2}},\\ H^{\alpha\beta}=\sqrt{-det(h)}h^{\alpha\beta}.\\$ (28)
Taking advantage of the membrane paradigm Iqbal and Liu (2009), as derived by the authors of Giataganas and Soltanpanahi (2014a, b), one can directly reads LGV-coefficients without solving motion equation,
@@
We rederived the formulas for our Langevin diffusion coefficients Eq. (29), Eq. (30), as well as the crucial intermediate formulas Eq. (24) and Eq. (26), by following Giataganas and Soltanpanahi (2014a, b) where the authors explained that these formulas are clear and explicit.
@@
[…]
@@
[…]
@@ III.1 Directions dependence
Initially, in this subsection, we will examine two extreme cases, in the ultra fast limit and at a fixed sufficiently small speed, to study the relations between different LGV-coefficients in all scale with $\mathcal{B}/T^{2}$, which is the only dimensionless parameter in this model. The numerical results of transverse LGV-coefficients $\kappa_{\perp}$ and longitudinal LGV-coefficients $\kappa_{\parallel}$ as a function of magnetic field $B/T^{2}$ are displayed in Fig. 2 and Fig. 3, normalized by the conformal limit given in (40). Fig. 2 shows transverse LGV-coefficients $\kappa^{v\parallel B}_{\perp}$, $\kappa^{v\perp B}_{(\perp,\parallel)}$ and $\kappa^{v\perp B}_{(\perp,\perp)}$, while Fig. 3 show longitudinal LGV-coefficients $\kappa^{v\parallel B}_{\parallel}$ and $\kappa^{v\perp B}_{\parallel}$.
When moving at the ultra fast limit, as shown in plot (a) in Fig. 2, the transverse LGV-coefficients of quark $\kappa_{\perp}$ has a relation reads $\kappa_{\perp}^{v\parallel B}\leq\kappa^{v\perp B}_{(\perp,\parallel)}\leq\kappa^{v\perp B}_{(\perp,\perp)}$ and $\kappa^{v\perp B}_{(\perp,\parallel)}\sim\kappa^{v\perp B}_{(\perp,\perp)}$, which means the transverse LGV-coefficients of quark mostly depending on the direction of moving but not the direction of transverse momentum diffusion when the moving velocity approaching to light speed. Since $\kappa^{v\perp B}_{(\perp,\parallel)}$ and $\kappa^{v\perp B}_{(\perp,\perp)}$ share the same motion direction, perpendicular to the magnetic field. It also means that the anisotropy induced by a uniform magnetic field in plane xozxoz contributes little to the strength of transverse LGV-coefficients when moving in the ultra-fast limit. But when the moving velocity is sufficiently small, $v=0.1$ is taken for example, one will find an opposite conclusion that the transverse LGV-coefficients of quark depend more on the direction of transverse momentum diffusion rather than the direction of moving. Since one can find $\kappa^{v\perp B}_{(\perp,\perp)}\sim\kappa^{v\parallel B}_{(\perp)}$ on plot (b) in Fig.(2) and they share the same transverse momentum diffusion direction perpendicular to the magnetic field (We will further examine this conclusion in the next subsection).
As for the longitudinal LGV-coefficients shown in Fig. 3, we consistently find that $\kappa_{\parallel}^{v\perp B}>\kappa_{\parallel}^{v\parallel B}$ both in ultra fast limit and at a fixed sufficiently small speed, which means moving perpendicular to $\mathcal{B}$ direction affects the longitudinal LGV-coefficients more strongly compared to motion along $\mathcal{B}$ direction. By comparing transverse LGV-coefficients in Fig. 2 with the longitudinal LGV-coefficients in Fig. 3, we find that the transverse LGV-coefficients are more sensitive to the moving velocity in the anisotropic magnetic background than the longitudinal LGV-coefficients.
@@
III.2 Velocity dependence
We further examine the impact of moving velocity $v$ on both transverse and longitudinal LGV-coefficients. In Fig. 4, we plot the LGV-coefficients as a function of quark velocity $v$, accompanied by the corresponding results in the conformal limit given in (40). It demonstrates that the increasing moving velocity $v$ enhances the LGV-coefficients in all cases we mentioned in this anisotropic magnetic background, which is very similar to the results in the conformal limit. This conclusion is also similar to the outcomes in spatially anisotropic case in Giataganas and Soltanpanahi (2014b).
Unlike the two extreme cases with fixed speeds in subsection III.1, we can study the same question with varying moving speeds. It is found that, at a sufficiently small moving speed, the transverse momentum broadening of a heavy quark depends more on the direction of momentum diffusion rather than the direction of moving with both $\kappa^{v\perp B}_{(\perp,\parallel)}<\kappa^{v\parallel B}_{\perp}\leq\kappa^{v\perp B}_{(\perp,\perp)}$ and $\kappa^{v\perp B}_{(\perp,\perp)}\sim\kappa^{v\parallel B}_{\perp}$ in the upper panel of Fig. (4). ($\kappa^{v\perp B}_{(\perp,\perp)}$ and $\kappa^{v\parallel B}_{\perp}$ share the same transverse diffusion direction, perpendicular to the magnetic field.) It is also found that the longitudinal LGV-coefficients, shown in the lower panel of Fig. 4, $\kappa_{\parallel}^{v\perp B}>\kappa_{\parallel}^{v\parallel B}$ at all fixed speeds.